Circular No. 8893

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2007 V1 (LARSON) S. M. Larson reports his discovery of a comet (discovery observation

tabulated below) on CCD images obtained with the 0.68-m Schmidt reflector in the course of the Catalina Sky Survey; he reported a 40′′ tail in p.a. 230^{o}. Following posting on the Minor Planet Center’s NEOCP webpage, several other CCD astrometrists have commented on the comet’s physical appearance. P. Birtwhistle (Great Shefford, Berkshire, England, 0.40-m f/6 Schmidt-Cassegrain reflector, Nov. 9.0 UT) found a concentrated coma of diameter 6′′ with a 30′′ tail in p.a. 235^{o}. J. Lacruz (Madrid, Spain, 0.40-m Ritchey-Chrétien reflector, Nov. 9.0) measured a 30′′ tail in p.a. 240^{o}. G. Hug (Eskridge, KS, U.S.A., 0.7-m reflector, Nov. 9.4) reports detecting “a short defined tail” in p.a. ∼ 250^{o}. E. Guido and G. Sostero (Castellammare di Stabia, Italy, remotely using a 0.25-m f/3.4 reflector located near Mayhill, NM, U.S.A., Nov. 9.4) write that a co-addition of forty 60-s unfiltered exposures on Nov. 9.4 reveals the presence of a diffuse coma nearly 12′′ in diameter with total mag ∼ 18.3. J. G. Ries (McDonald Observatory, 0.76-m reflector + prime-focus corrector, Nov. 9.3) finds a very clear tail ≈ 64′′ long pointing almost exactly southwest. R. E. Hill (Mt. Lemmon 1.5-m reflector, Nov. 9.4) adds that a co-addition of four 60-s exposures shows a broad 7′ tail in p.a. 240^{o} and a well-defined nuclear condensation 5′′–6′′ across; the total mag was given as 17.7.

2007 UT  α_{2000}  δ_{2000}  Mag.
Nov. 8.31007  3^{h}14^{m}10.^{s}71  +30^{o}01′12′′.8  17.3

The available astrometry, the following elliptical orbital elements, and an ephemeris appear on MPEC 2007-V72.

T = 2007 Dec. 8.5940 TT  ω = 51.^{o}4714
e = 0.461170  Ω = 8.1899 }2000.0
q = 2.676678 AU  i = 10.7903
a = 4.967569 AU n^{o} = 0.0890201 P = 11.07 years

COMETS C/2004 E5 (SOHO) Further to IAUC 8892, another slightly diffuse Kreutz-sungrazing com-

et of mag ∼ 7.5 has been found on archival SOHO website images.

Comet 2004 UT α_{2000} δ_{2000} Inst. F MPEC
C/2004 E5 Mar. 1.321 23^{h}03.^{m}3 − 8^{o}08′ C3 BZ 2007-U23
2007 November 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green